An athlete takes 2 sec to reach his max speed of 18km/h. What is the magnitude of his average acceleration?

Grade:12th pass

1 Answers

Arun
6 years ago
initial velocity= 0
final velocity= 8km/h
time= 2s units of time and velocity are different, first make them equal.
8km/h= 8*1000
              1*3600
          =8000 
            3600
 
          =5m/s
a= v-u/t
  = 5-0
       2
   =5/2 =2.5m/s2